OpenStudy (anonymous):

MEDAL!!! Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2.4 h, and Car B traveled the distance in 4 h. Car A traveled 22 mph faster than Car B. How fast did Car A travel?

ganeshie8 (ganeshie8):

do you know the relation between distance, time and speed ?

OpenStudy (anonymous):

Uhhmm no

ganeshie8 (ganeshie8):

distance = speed x time

OpenStudy (anonymous):

Ok can u show me how to get the answer plz

ganeshie8 (ganeshie8):

how much `distance` you had travelled depends on `speed` at which you travelled and the `time` you took to travel

ganeshie8 (ganeshie8):

Say speed of Car A is `s`

OpenStudy (anonymous):

Ok

ganeshie8 (ganeshie8):

` Car A traveled 22 mph faster than Car B` that means speed of carB would be `s-22`, yes ?

OpenStudy (anonymous):

I'm guessing so

ganeshie8 (ganeshie8):

Car A : speed = s time = 2.4 distance = s*2.4 Car B : speed = s-22 time = 4 distance = (s-22)*4

ganeshie8 (ganeshie8):

since both cars covered the same distance, set the distances equal to each other and solve s

ganeshie8 (ganeshie8):

s*2.4 = (s-22)*4 s = ?

OpenStudy (anonymous):

It woud be

ganeshie8 (ganeshie8):

distribute on the right hand side and isolate s

OpenStudy (anonymous):

Oh wait my computer is freazzinf

OpenStudy (anonymous):

Let try and figure it out real quick

ganeshie8 (ganeshie8):

take your time :)

ganeshie8 (ganeshie8):

s*2.4 = (s-22)*4 distribute on the right hand side, what do you get ?

OpenStudy (anonymous):

Is the (*) sign exponents?

ganeshie8 (ganeshie8):

Oh, * is multiplication

OpenStudy (anonymous):

Just making sure(:

OpenStudy (anonymous):

S=55?

ganeshie8 (ganeshie8):

Yes!

OpenStudy (anonymous):

Thx((:

ganeshie8 (ganeshie8):

yw!

OpenStudy (anonymous):

\(\Huge\cal\color{blue}{your~very~good~at~answering~things}\)